Please help me find the area of a triangle if:
a=2, b=1, and C=100 degrees.
The choices are as follows:
1. 0.98
2. 1.97
3. 0.17
4. 1.53
Thanks, Tami
: Please help me find the area of a triangle if:
a=2, b=1, and C=100 degrees.
The choices are as follows:
1. 0.98
2. 1.97
3. 0.17
4. 1.53
Thanks, Tami

Answer by josmiceli(2023)

The angle is 100 degrees which is 10 degrees more than a
right angle. The formula for area is (1/2)*b*h,
where b is any of the sides, and once you've
chosen that side for a base, then you find the height, h
I'll choose the side = 1 for a base. The angle, C = 100degrees
is between this base and the side that = 2
I'll drop a vertical from the end of the side = 2, and that
will be my height
Have you had trig, or is this a trig question?
If so, sin (80) = h/2
.9848 = h/2
h = 1.9696
(1/2)*b*h = (1/2)*1*1.9696
A = .9848 answer
It looks like the 1st answer is correct